Based on below data, test that Brand preference and Education qualifications are independent, at 5% level of significance.


(No. OF PERSONS PREFERRING BRANDS)


Education Brand A Brand B Brand C

qualification


SSC 22 34 35

HSC 28 35 28

Graduate 30 44 27

Post graduate 55 42 28


1
Expert's answer
2021-02-24T06:43:56-0500

The Chi-Square Test of Independence:

"H_0" : Brand preference is independent of Education qualifications

"H_1" : Brand preference is dependent of Education qualifications


"\Chi^2=\displaystyle\sum^R_{i=1}\displaystyle\sum^C_{j=1}\frac{(a_{ij}-e_{ij})^2}{e_{ij}}"


"e_{ij}=\frac{row(i)total\cdot col(j)total}{grandtotal}"


"grandtotal=408"


"e_{11}=\frac{135\cdot91}{408}=30.11" , "e_{12}=\frac{155\cdot91}{408}=34.57" , "e_{13}=\frac{118\cdot91}{408}=26.32"

"e_{21}=\frac{135\cdot91}{408}=30.11" , "e_{22}=\frac{155\cdot91}{408}=34.57" , "e_{23}=\frac{118\cdot91}{408}=26.32"

"e_{31}=\frac{135\cdot101}{408}=33.42" , "e_{32}=\frac{155\cdot101}{408}=38.37" , "e_{33}=\frac{118\cdot101}{408}=29.21"

"e_{41}=\frac{135\cdot125}{408}=41.36" , "e_{42}=\frac{155\cdot125}{408}=47.49" , "e_{43}=\frac{118\cdot125}{408}=36.15"


"\Chi^2=2.18+0.01+2.86+0.15+0.01+0.11+0.35+0.83+0.17+"

"+4.50+0.63+1.84=13.64"


Degrees of freedom:

"df=" (number of columns – 1)(number of rows – 1)

"df=2\cdot3=6"


p-value"=P(\Chi^>13.64)=0.038"

"\alpha=0.05" , "\alpha>" p-value

Since α > p-value, reject H0. This means that the factors are not independent.